# Variational Scarring in Graphene Quantum Dots

### v2

@@ Abstract
In quantum scarring, the probability density of a quantum state is imprinted with a short, unstable periodic orbit (PO) of a chaotic classical system due to quantum interference. These scars are surprising as there is no such classical concentration of density near a PO in the long time limit for a bounded chaotic system, hence standing out against the corresponding, monotonous backdrop of quantum state randomness across most of phase space. Since the initial prediction of the phenomenon by Heller [9], extensive experimental attempts have been carried out to image scarred states. Even though analogues of scars have been observed in various classical wave experiments, such as in microwave resonators [12, 13], acoustic cavities [14], and fluid surface waves [15], their direct visualization in a genuine quantum system remained elusive until the recent breakthrough  [16] in a stadium-shaped, graphene-based quantum dot (QD).
@@
Inspired by this progress, we suggest an experimental pathway to directly visualize a different recently discovered member of the scar family, namely variational scars, first reported in disordered two-dimensional nanostructures [17]. Although these scars resemble the conventional scarring by Heller [9], they stem from a fundamentally different mechanism [17, 18]: the (near) symmetry of the unperturbed system leads to scarred states favored by the variational principle in the presence of a localized perturbation. As a result, the scars in the perturbed quantum system form around the POs of the unperturbed classical counterpart, so we call them perturbation-induced, variational scars. This scarring mechanism is ubiquitous [17, 18, 19, 20, 21, 22, 23, 24], but experimental validation of variational scarring has yet to be achieved. To address this issue, we investigate the potential of a realizable quantum dot on monolayer graphene to foster variational scars, allowing a STM detection of these scarred states, in the same spirit as in Ref. [16].
@@
Figure 2: The figure summarizes the observed features of the variational scars in the elliptical QD perturbed by a nanotip. Subplot (a) displays a collection of high-energy eigenstates of the Hamiltonian described by Eq. 1, with some strongly scarred by classical Lissajous periodic orbits. Subplot (b) highlights an exemplary scar, where the probability density concentrates along the Lissajous orbit in the shape of the symbol alpha (blue curve). The scar is also oriented in such a way that its loop avoids the bump (violet dot), effectively minimizing the perturbation caused by the nanotip. These variational scars are accompanied by companion scars that, as shown in subplot (c), maximize the the effect of perturbation, rather than minimize it. Some of the displayed quantum states exhibit classical bouncing-ball-like motion (blue dashed circle), while others remain almost unaffected by the nanotip (red dashed circle).
@@
$V_{\textrm{ext}}(\mathbf{r})=\frac{1}{2}\omega_{0}\left(\alpha^{2}x^{2}+y^{2}\right)$ (2)
where $\omega_{0}$ is the overall confinement strength and $\alpha$ represents the ellipticity. We set $\omega_{0}=2$, aligning with the experimental confinement strengths reported for similar QDs. Without loss of generality, we focus on a special value of $\alpha=0.5$ where, crucially, accidental degeneracies occur in the unperturbed system associated with the Lissajous orbits existing in the classical counterpart. We note that Lissajous scars appear at various values of ellipticity $\alpha$ (see Ref. [18]). Moreover, as will be shown later, the ellipticity $\alpha$ does not need to be precisely matched to the classical resonant conditions of the Lissajous POs.
@@
described by its amplitude $A=0.35$ and width $\sigma=1.8$. Similarly, randomly distributed Gaussian bumps are also employed to model impurity noise in a QD [35, 36, 37]. Here, we have selected the values of the amplitude and width so that the nanotip induces strong scarring in the relevant energy regime. Although we spotlight the local perturbation caused by a nanotip in this study, a similar effect can be engineered by introducing an adatom(s) at a specific location(s) on the graphene layer, or by leaving an n-doped region within the p-doped domain that forms the QD during the manufacturing phase. These alternatives avoid the difficulty of arranging an STM tip in conjunction with the nanotip that generates the perturbation $V_{\textrm{tip}}$. Nonetheless, a multi-tip approach would provide a new level of controllability (see, e.g. Refs. [38, 39, 40]).
@@
We emphasize that this type of scarred state does not exist in the unperturbed QD, whose eigenstates $|\phi_{n}\rangle$ are instead expressible in terms of Hermitian polynomials  [^2] The eigenstates of an elliptical are $|\phi_{n,m}\rangle\propto H_{m}(\sqrt{\alpha\omega_{0}}x)H_{n}(\sqrt{\omega_{0}}y)\exp[-\omega_{0}(\alpha^{2}x^{2}+y^{2})/2]$, where $H_{m}(\cdot)$ is the Hermite polynomial of order $m$. The corresponding energy spectrum shows degeneracies at commensurable frequencies, that is, when the ellipticity is a rational number.. This observation strongly supports the assertion that the scars in the QD are variational in nature. To further confirm the origin of the reported scarring, we show two other trademarks unique to variational scarring. First, some of the observed scars align themselves to minimize the perturbation caused by the nanotip, as seen in Fig. 2 (b); whereas the others counterintuivitely seek to maximize it, as displayed in Fig. 2 (c). Second, we expand the scar presented in Fig. 2 (b) in terms of the eigenstates of the unperturbed elliptical QD as $|\textrm{Scar}\rangle=\sum_{n}c_{n}|\phi_{n}\rangle$. This type of “subspectrum” analysis of the expansion coefficients $|c_{n}|^{2}$ reveals that the scars are formed by a small set of states close in energy, which is consistent with the theory of variational scarring [17, 20, 19, 18, 22]. For instance, the showcased Lissajous scar is primarily ($\sim 90\,\%$) conjured by three nearly-degenerate eigenstates of the elliptical QD without the perturbation, as shown in Fig. 3.
@@
In contrast to the typical scenario involving multiple localized perturbations, we can selectively target scars of different geometries corresponding to distinct classical resonances based on the position of the nanotip. For instance, instead of the $(2,3)$ resonance scar shown in Figure 2(b) resembling the alpha shape, the ellipticity $\alpha=0.5$ would typically produce scars from the $(1,2)$ resonance, manifesting as either a parabola or a figure-eight. Both resonances are energetically close to each other. More explicitly, the $(1,2)$ resonance exhibits exact degeneracy, while the $(2,3)$ resonance is only approximately degenerate for $\alpha=0.5$. However, unlike the case of many random bumps [18], the scarred state from the $(2,3)$ resonance offers more effective extremization of the perturbation due to the precise placement of the nanotip. Notably, the self-crossing point of the $(2,3)$ resonance provides a particular advantage in minimizing and maximizing the perturbation (see, Fig. 2(b) and (c), respectively).
On the other hand, moving the nanotip affects the level of scarring, as illustrated in Fig. 4 (c), for three different tip positions shifted from the center of the QD by $\Delta x=2,6,10$ along the positive x-axis [the “optimized” scars shown in Fig. 2(b) and (c) have $\Delta x=10$]. The scars are weaker when the bump caused by the tip does not align with a position where the (1,2) resonance scar can optimally minimize or maximize the perturbation, i.e., by coinciding with or avoiding the self-crossing point of the scar and the bump. Nevertheless, noticeable concentrations of probability density along the corresponding Lissajous orbits are observed in some of the perturbed states. Furthermore, since the potential and, consequently, the periodic orbits scale with energy, the effect of the tip position can be mitigated by either focusing on a different energy range or slightly adjusting the confining potential.
@@
In particular, we highlight a situation where the probing bias $V_{s}$ is adjusted to match with the scarred state of Fig. 2(b). Figure 5 (a) presents the corresponding spatial distribution of the LDOS that paints a discernible pattern resembling the target scar at moderate energy resolution $\eta=10^{-4}$ with $\Delta V=0$. Remarkably, the scar pattern remains apparent even when the probe voltage is mismatched, like in Fig. 5(b) with $\Delta V=3.5\cdot 10^{-4}$, or when the energy resolution is degraded to $\eta=10^{-3}$, as shown in Fig. 5(c). This simple example thus underlines the point that the variational scarring can indeed lead to a resilient observable STM signature. In other words, the STM can provide an experimentally feasible way to directly visualize the effect of variational scars, as opposed to the prevailing approach of deducing their presence incidentally through transport measurements, such as in Refs. [52, 53, 54, 55, 56].
@@
Figure 5: Figure shows the constructed LDOS map for a hypothetical differential conductance $\partial I/\partial V$ measurement. The probing bias $V_{s}$ of the simulated STM tip is adjusted to match with the example scar of Fig. 2 (b). As seen in (a), the corresponding LDOS map contains a distinct imprint of the Lissajous orbit associated with the targeted scarred state when $\Delta V=0$ and $\eta=10^{-4}$. The scar pattern persists even when $\Delta V=3.5\cdot 10^{-3}$ in (b) and $\eta=10^{-3}$ in (c).
@@
In addition to facilitating the observation of variational scars, the proposed approach may be also employed to explore a necessary but subtle consequence of quantum scarring, namely antiscarring [57, 58]. An interesting avenue of future research is to investigate if there exists a hybrid form of variational [17, 20, 19, 18] and relativistic scars [59, 60, 61], or a possible transition between these two types of scarring, as the energy is varied. Another interesting possibility is to utilize the presented quantum device as a platform to study quasi-scarred states in a spirit similar to that of Ref. [62] that might occur due to the roughness and atomic structure of the underlying monolayer graphene, and lattice scars within a more general framework in contrast to the previous study [63].
@@
In summary, we have demonstrated the existence of variational scars induced by a nanotip in a graphene QD. Furthermore, the effect of perturbation-induced scarring has been shown to be detectable in STM measurements, marking a significant step towards directly visualizing variational scars.
@@ Acknowledgements.
We are thankful for the useful discussions with E. Räsänen, A. Mert Bozkur, and Z. Ge. Furthermore, J.K.-R. thanks the Oskar Huttunen Foundation for financial support. This project is also supported by the National Science Foundation (Grant No. 2403491).

### v3

@@ Abstract
In quantum scarring, the probability density of a quantum state is imprinted with a short, unstable periodic orbit (PO) of a chaotic classical system due to quantum interference. These scars are surprising as there is no such classical concentration of density near a PO in the long time limit for a bounded chaotic system, hence standing out against the corresponding, monotonous backdrop of quantum state randomness across most of phase space. Since the initial prediction of the phenomenon by Heller [9], extensive experimental attempts have been carried out to image scarred states. Even though analogs of scars have been observed in various classical wave experiments, such as in microwave resonators [12, 13], acoustic cavities [14], and fluid surface waves [15], their direct visualization in a genuine quantum system remained elusive until the recent breakthrough  [16] in a stadium-shaped, graphene-based quantum dot (QD).
@@
Inspired by this progress, we suggest an experimental pathway to directly visualize a different recently discovered member of the scar family, namely variational scars, first reported in disordered two-dimensional nanostructures [17]. Although these scars resemble the conventional scarring by Heller [9], they stem from a fundamentally different mechanism [17, 18]: the (near) symmetry of the unperturbed system leads to scarred states favored by the variational principle in the presence of a localized perturbation. As a result, the scars in the perturbed quantum system form around the POs of the unperturbed classical counterpart, so we call them perturbation-induced, variational scars. This scarring mechanism is ubiquitous [17, 18, 19, 20, 21, 22, 23, 24], but experimental validation of variational scarring has yet to be achieved. To address this issue, we investigate the potential of a realizable quantum dot on monolayer graphene to foster variational scars, allowing a scanning tunneling microscopy (STM) detection of these scarred states, in the same spirit as in Ref. [16].
@@
Figure 2: The figure summarizes the observed features of the variational scars in the elliptical QD perturbed by a nanotip. Subplot (a) displays a collection of high-energy eigenstates of the Hamiltonian described by Eq. 1, with some strongly scarred by classical Lissajous periodic orbits. Subplot (b) highlights an exemplary scar [marked by the red square in (a)], where the probability density concentrates along the Lissajous orbit in the shape of the symbol alpha (blue curve). The scar is also oriented in such a way that its loop avoids the bump (violet dot), effectively minimizing the perturbation caused by the nanotip. These variational scars are accompanied by companion scars that, as shown in subplot (c), maximize the the effect of perturbation, rather than minimize it. Some of the displayed quantum states exhibit classical bouncing-ball-like motion (blue dashed circle), while others remain almost unaffected by the nanotip (green dashed circle).
@@
$V_{\textrm{ext}}(\mathbf{r})=\frac{1}{2}\omega_{0}\left(\alpha^{2}x^{2}+y^{2}\right),$ (2)
where $\omega_{0}$ is the overall confinement strength and $\alpha$ represents the ellipticity. We set $\omega_{0}=2$, aligning with the experimental confinement strengths reported for similar QDs. Without loss of generality, we focus on a special value of $\alpha=0.5$ where, crucially, accidental degeneracies occur in the unperturbed system associated with the Lissajous orbits existing in the classical counterpart. We note that Lissajous scars appear at various values of ellipticity $\alpha$ (see Ref. [18]). Moreover, as will be shown later, the ellipticity $\alpha$ does not need to be precisely matched to the classical resonant conditions of the Lissajous POs. Finally, we emphasize that the experimental feasibility of this type of graphene-based quantum dot has been substantiated by recent works, such as Refs. [16, 26, 25].
@@
described by its amplitude $A=0.35$ and width $\sigma=1.8$. Similarly, randomly distributed Gaussian bumps are also employed to model impurity noise in a QD [36, 37, 38]. Here, we have selected the values of the amplitude and width so that the nanotip induces strong scarring in the relevant energy regime. Although we spotlight the local perturbation caused by a nanotip in this study, a similar effect can be engineered by introducing an adatom(s) at a specific location(s) on the graphene layer, or by leaving an n-doped region within the p-doped domain that forms the QD during the manufacturing phase. These alternatives avoid the difficulty of arranging an STM tip in conjunction with the nanotip that generates the perturbation $V_{\textrm{tip}}$. Nonetheless, a multitip approach would provide a new level of controllability (see, e.g. Refs. [39, 40, 41]).
@@
We emphasize that this type of scarred state does not exist in the unperturbed QD, whose eigenstates $|\phi_{n}\rangle$ are instead expressible in terms of Hermitian polynomials  [^3] The eigenstates of an elliptical are $|\phi_{n,m}\rangle\propto H_{m}(\sqrt{\alpha\omega_{0}}x)H_{n}(\sqrt{\omega_{0}}y)\exp[-\omega_{0}(\alpha^{2}x^{2}+y^{2})/2]$, where $H_{m}(\cdot)$ is the Hermite polynomial of order $m$. The corresponding energy spectrum shows degeneracies at commensurable frequencies, that is, when the ellipticity is a rational number.. This observation strongly supports the assertion that the scars in the QD are variational in nature. To further confirm the origin of the reported scarring, we show two other trademarks unique to variational scarring. First, some of the observed scars align themselves to minimize the perturbation caused by the nanotip, as seen in Fig. 2 (b); whereas the others counterintuivitely seek to maximize it, as displayed in Fig. 2 (c). Second, we expand the scar presented in Fig. 2 (b) in terms of the eigenstates of the unperturbed elliptical QD as $|\textrm{Scar}\rangle=\sum_{n}c_{n}|\phi_{n}\rangle$. This type of “subspectrum” analysis of the expansion coefficients $|c_{n}|^{2}$ reveals that the scars are formed by a small set of states close in energy, which is consistent with the theory of variational scarring [17, 20, 19, 18, 22]. For instance, the showcased Lissajous scar is primarily ($\sim 90\,\%$) conjured by three nearly degenerate eigenstates of the elliptical QD without the perturbation, as shown in Fig. 3.
@@
In contrast to the typical scenario involving multiple localized perturbations [19, 20, 18, 21, 17], we can selectively target scars of different geometries corresponding to distinct classical resonances based on the position of the nanotip. For instance, instead of the $(2,3)$ resonance scar shown in Fig. 2(b) resembling the alpha shape, the ellipticity $\alpha=0.5$ would typically produce scars from the $(1,2)$ resonance, manifesting as either a parabola or a figure-eight. Both resonances are energetically close to each other. More explicitly, the $(1,2)$ resonance exhibits exact degeneracy, while the $(2,3)$ resonance is only approximately degenerate for $\alpha=0.5$. However, unlike the case of many random bumps [18], the scarred state from the $(2,3)$ resonance offers more effective extremization of the perturbation due to the precise placement of the nanotip. Notably, the self-crossing point of the $(2,3)$ resonance provides a particular advantage in minimizing and maximizing the perturbation [see, Fig. 2(b) and (c), respectively].
On the other hand, moving the nanotip affects the level of scarring, as illustrated in Fig. 4 (c), for three different tip positions shifted from the center of the QD by $\Delta x=2,6,10$ along the positive x axis [the “optimized” scars shown in Figs. 2(b) and (c) have $\Delta x=10$]. The scars are weaker when the bump caused by the tip does not align with a position where the (1,2) resonance scar can optimally minimize or maximize the perturbation, i.e., by coinciding with or avoiding the self-crossing point of the scar and the bump. Nevertheless, noticeable concentrations of probability density along the corresponding Lissajous orbits are observed in some of the perturbed states. Furthermore, since the potential and, consequently, the periodic orbits scale with energy, the effect of the tip position can be mitigated by either focusing on a different energy range or slightly adjusting the confining potential.
@@
In particular, we highlight a situation where the probing bias $V_{s}$ is adjusted to match with the scarred state of Fig. 2(b). Figure 5 (a) presents the corresponding spatial distribution of the LDOS that paints a discernible pattern resembling the target scar at moderate energy resolution $\eta=10^{-4}$ with $\Delta V=0$. Remarkably, the scar pattern remains apparent even when the probe voltage is mismatched, like in Fig. 5(b) with $\Delta V=3.5\times 10^{-4}$, or when the energy resolution is degraded to $\eta=10^{-3}$, as shown in Fig. 5(c). Notably, the considered resolution range, $10^{-3}\leq\eta\leq 10^{-4}$, lies well within current experimental capabilities, as demonstrated, for example, in Ref. [16].
Whereas previous works, such as, Refs. [19, 20, 18, 17] have primarily focused on establishing the theoretical foundation for variational scarring; our work advances this foundation towards concrete experimental realization, within a highly tunable platform such of graphene exhibiting unique features, such as pseudo-relativistic behavior [16, 25, 26]. The simple example above thus underlines the point that the variational scarring can indeed lead to a resilient observable STM signature. In other words, the STM can provide an experimentally feasible way to directly visualize the effect of variational scars, as opposed to the prevailing approach of deducing their presence incidentally through transport measurements, such as in Refs. [53, 54, 55, 56, 57]. Furthermore, we note that previously considered variational scarring within a semiconductor QDs, such as studied in Refs. [19, 20, 18, 22], is not well-suited for an STM observation, thereby enhancing the appeal of the strategy proposed here. While optical analogs, such as the one presented in Ref. [18], provide an alluring opportunity, the suggested STM scheme offers significant value in observing variational scarring within a genuine quantum environment.
@@
Figure 5: Figure shows the constructed LDOS map for a hypothetical differential conductance $\partial I/\partial V$ measurement. The probing bias $V_{s}$ of the simulated STM tip is adjusted to match with the example scar of Fig. 2 (b). As seen in (a), the corresponding LDOS map contains a distinct imprint of the Lissajous orbit associated with the targeted scarred state when $\Delta V=0$ and $\eta=10^{-4}$. The scar pattern persists even when $\Delta V=3.5\times 10^{-3}$ in (b) and $\eta=10^{-3}$ in (c).
@@
In addition to facilitating the observation of variational scars, the proposed approach may be also employed to explore a necessary but subtle consequence of quantum scarring, namely antiscarring [58, 59]. An interesting avenue of future research is to investigate if there exists a hybrid form of variational [17, 20, 19, 18] and relativistic scars [60, 61, 62], or a possible transition between these two types of scarring, as the energy is varied. Another interesting possibility is to utilize the presented quantum device as a platform to study quasiscarred states in a spirit similar to that of Ref. [63] that might occur due to the roughness and atomic structure of the underlying monolayer graphene, and lattice scars within a more general framework in contrast to the previous study [64].
@@
In summary, we have demonstrated the existence of variational scars induced by a nanotip in a graphene QD, extending the validity of the scar theory. Furthermore, the effect of perturbation-induced scarring has been shown to be detectable in STM measurements, marking a significant step towards directly visualizing variational scars very first time.
@@ Acknowledgements.
We are thankful for the useful discussions with E. Räsänen, A. Mert Bozkur, and Z. Ge. Q. Y., T. Z, and J.V.J., acknowledge support from the Gordon and Betty Moore Foundation award 10.37807/GBMF11569. Furthermore, J.K.-R. thanks the Oskar Huttunen Foundation for financial support. This project is also supported by the National Science Foundation (Grant No. 2403491).